OpenStudy (anonymous):

If A and B are independent events, P(A)=.9, and P(B)=.7, find P(A U B)

OpenStudy (anonymous):

\[P(A\cup B)=P(A)+P(B)-P(A\cap B)\] Since \(A\) and \(B\) are independent, this means \(P(A\cap B)=P(A)P(B)\). \[P(A\cup B)=P(A)+P(B)-P(A)P(B)\]

OpenStudy (anonymous):

.97 ??

OpenStudy (zarkon):

yes